# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

### v3

@@ II.3 Transport dynamics
Under the Born-Markov approximations and the rotating-wave approximation, the dynamics of the system that exists dissipation at each TLS in the multichains, the coupling between the first site of the first chain and the driving field as well as between the last site of each chain and the reservoir, are governed by the Lindblad quantum master equation lindblad1976generators . Then the exciton transport dynamics in the system can be investigated by the master equation
@@
where the first term in the righthand side accounts for the free evolution under the total Hamiltonian (11), the second term characterizes the decay of excitons due to the spontaneous emission with
$D_{{\rm d},k}[\rho]=\sum^{N}_{i=1}\gamma\mathcal{L}_{\sigma^{-}_{k,i}}[\rho]+\kappa\mathcal{L}_{\rm c}[\rho],$ (14)
where $\gamma$ and $\kappa$ are the decay rates of the TLSs and the cavity photons respectively, and $\mathcal{L}_{x}[\rho]=x\rho x^{\dagger}-\dfrac{1}{2}\left\{\rho,x^{\dagger}x\right\}$ is the Lindblad superoperators with standard form. The term $D_{{\rm R},k}[\rho(t)]$ in Eq. (13) represents the exciton exchange between the last site of the $k$th chain and the exciton drain, which is given by
@@
where $\gamma_{\rm R}$ is the exciton exchanging rate between the last site of each chain and the drain, $\bar{n}_{\rm R}$ is the average exciton number in the drain. The last sites of each chain exist coupling with the drain, so $k\in\{1,2,...,L\}$. In Eq. (15), the two terms denote respectively the inflow and outflow of excitons from the last column TLSs of the multichains to the drain. In order to show the net transportation of exciton, we set as $\bar{n}_{\rm R}=0$ which is reasonable as the photosynthesis process in nature. Therefore, the exciton is unidirectionally transferred from the last column of the multichains to the drain without exciton reflection.
@@ II.4 Efficiency of exciton transport
Figure 2: (color online) The energy spectra of the exciton-cavity coupling system as a function of the inter-chain coupling strength $\zeta$ for different configurations HO ST (a), HE ST (b), HO-HE ST (c), with the colormap representing the Hopfield coefficients of the content of cavity photons by solving the eigenvalues and eigenstates of the Hamiltonian (1) in zero- and single-excitation subspace. (d-f) The second order partial derivative of the Hopfield coefficients of the content of cavity photons with respect to $\zeta$, which contain the discontinuity and the extremum corresponding to the marked locations existing evident changes in steady-state dynamics. (g-i) The corresponding von Neumann entropy $S$ of excitons by calculating Eq. (12). The color of the lines has not means just to distinguish each other in (d-i). The cavity-dressed energy level crossings (anticrossings) producing obvious effect on exciton transport are marked by the black (violet) vertical lines and the data texts of their locations, which correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy. The corresponding parameters are $L=2$, $N=8$, $\Delta_{\rm a}=\Delta_{\rm c}=0$ eV, $h=1$ eV, $\delta=0.5$, and $\Omega_{R}=1$ eV.
@@
To show the exciton transport dynamics of the system containing a 2D multichains and a cavity, the exciton transport efficiency is defined and derived in this subsection. Generally, the exciton transport efficiency is characterized by the ratio of the outgoing $(I^{\rm s}_{\rm o})$ and input $(I^{\rm s}_{\rm i})$ exciton currents for steady state liu2019vibration ; wei2019enhanced , that is
@@
where the superscript ”s” refers to the steady state. On the other hand, the exciton occupation number can display the transport machanism and be used to solve the $I_{\rm i}$ and $I_{\rm o}$. The exciton occupation number on the $j$th site of the $k$th chain is given by
@@
Then combining Eq. (13) with Eq. (17), the equation that $p_{k,j}(t)$ obeys is derived as
@@
where the trace is operated throughout the Hilbert space of the total system. The lefthand side of Eq. (18) means the rate of the change of $p_{k,j}(t)$ with time, thus the first term (the second term) of the righthand side represents the input (outgoing) exciton current at the $j$th site in the $k$th chain. According to the explanation, under the TLS and TLS+cavity drive the input exciton current of the multichains can be written respectively as
@@
$I_{\rm o}(t)={\rm Tr}[\sum^{L}_{k=1}D_{{\rm R},k}[\rho(t)]\sigma_{k,N}^{+}\sigma_{k,N}^{-}],$ (22)
where each chain has exciton outgoing to the drain for all configurations, thus $k\in\{1,2,...,L\}$. Pluging the Eq. (15) into the Eq. (22), under the zero- and single- excitation assumption, the form of the outgoing exciton current is simplified as
@@
where $p_{k,N}$ is the occupation number on the last site of the $k$th chain.
@@ III Spectrum of coupling system
There are a large number of studies on the spectra of the SSH Hamiltonian for the open boundary conditions sirker2014boundary ; nie2020bandgap ; wei2022cavity ; cai2019observation . The topological phases of the SSH chain depend on the parity of the number of an 1D TLSs chain and the qubit-qubit coupling configuration. According to the Ref. cai2019observation , the dimerization parameter $\delta>0$ ($\delta<0$) corresponds to the topological winding number $\nu=1$ ($\nu=0$), i.e., the topologically nontrivial (trivial) regime. In Ref. sirker2014boundary , the spectra show that there is one zero-energy mode in whole range of $\delta$ for odd and are two localized nearly-zero-energy modes in the topologically nontrivial regime $\delta>0$ for even chain. The finite gap between the two edge modes leads to the formation of the hybrid edge modes that overlap with each other, generating the polaritonic states when the edge modes are dressed by cavity due to the coupling between the excitons and the photons. Comparing to the single chain system, the presence of the inter-chain coupling induces the extension of the spectra of the 2D multi-chain system. As we all known, the gap corresponds to the energy level anticrossings, at which the LZ transitions among the levels occur in multi-level LZ system, adding greatly to the complexity of the system dynamics hu2018effect . What’s more, the energy level crossings mean that the corresponding quantum states are degeneracy. In the exciton-cavity coupling system, at the crossings the excitations are coherently exchanged between the TLSs and the cavity. Therefore, the crossings and anticrossings of the edge modes dressed by the cavity can lead to the abundant exchange of excitation between the multichains and the cavity, increasing or decreasing polaritons. The forming polaritons act as intermediary in exciton transport, where the excitons are transported among site $(k,1)~{}\rightarrow$ cavity $\rightarrow$ site $(k,N)$ without experiencing the dissipation of the other TLSs on multichains wei2022cavity . Therefore, the polaritons lead to the exciton transport efficiency and currents increasing, which can be controlled by the crossings and anticrossings.
Figure 6: (color online) The dependence of the steady-state occupation numbers (a1 and b1) $p_{11}$ on site 11, (a2 and b2) $p_{\rm bulk}$ on bulk sites, (a3 and b3) input exciton current $I_{\rm i}$, (a4 and b4) outgoing exciton current $I_{\rm o}$, and (a5 and b5) transport efficiency $\eta$ on the inter-chain coupling strength $\zeta$ (a1-5) and the dimerization parameter $\delta$ (b1-5) for different configurations without cavity. In (b1-5) the solid and dash lines correspond respectively to $N=8$ and 9. The parameters used are $\zeta=0.1$ eV, $\xi_{{\rm T},1}=\xi_{\rm c}=0.1$ eV, $\gamma_{\rm d}=0.01$ eV, $\gamma_{\rm R}=0.1$ eV, and the others are the same as those in Fig. 2.
@@
In order to analyse the spectra of the coupling system consisting of the multichains and the cavity, its eigenvalues can be obtained by numerically diagonalizing the Hamiltonian (1). Further, the formation of polaritons can be investigated by calculating the Hopfield coefficients $|\alpha_{c}|^{2}$. When the $j$th normalized eigenvector is expanded as $|\psi_{j}\rangle=\Sigma_{k}^{L}\Sigma_{i}^{N}\alpha_{k,i}|k,i\rangle+\alpha_{c}|c\rangle$, $|\alpha_{k,i;c}|^{2}$ denotes Hopfield coefficient, which characterizes the corresponding hybridizing weight of different excitonic and photonic contents. The color maps Fig. 2 (a-c), 3 (a-c) and Fig. 4 (a-d), 5 (a-d) show respectively the spectra of the exciton-cavity coupling system as functions of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$ for different coupling configurations, and contain the photonic Hopfield coefficient $|\alpha_{c}|^{2}$. One can see that the zero-energy modes exist in the spectra for each situation. $|\alpha_{c}|^{2}\neq 0$ means that the energy level is dressed by the cavity mode. It is obvious that only the nearly-zero-energy modes are distinctly dressed by the cavity mode, generating corresponding polaritonic branches, because the high excited states are hardly dressed for the given coupling strength between excitons and photons. In addition, there are the crossings (anticrossings) in the spectra, in which the crossings (anticrossings) between the nearly-zero-energy modes dressed by the cavity mode are our focus. It can be seen that the ratio of the distributed photonic contents on the upper and lower polaritonic branches vary with $\zeta$ and $\delta$. It means that the coherent exchange of excitation or the LZ transition of photons occurs, inducing the change of the exciton transport efficiency and currents, which will be analysed in the next section by the steady-state dynamics. Further, it can be discovered that the number and locations of crossings (anticrossings) affecting transport depend on the coupling configurations of TLSs in the range of $\zeta$ and $\delta$. In comparison, the effect of $\zeta$ on the crossings (anticrossings) is more significant than that of $\delta$.
@@
In order to find the changing regularity of the photonic Hopfield coefficient $|\alpha_{c}|^{2}$, we plot the scatter maps of the second order partial derivatives $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ of $|\alpha_{c}|^{2}$ with respect to $\zeta$ in Fig. 2 (d-f), 3 (d-f) and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ with respect to $\delta$ in Fig. 4 (e-i), 5 (e-i) that are related to the obvious changes of the steady-state dynamics of exciton transport. To determine and distinguish accurately the locations of the crossings and anticrossings in spectra, the corresponding von Neumann entropy $S$ of excitons is calculated through Eq. (12), and is shown in Fig. 2 (g-i), 3 (g-i), 4 (i-l), and 5 (i-l). The crossings (anticrossings) determined by $S$ is signed in the figure of the spectra, which correspond to the saltatorial intersections (maximum values or acclive intersections) of $S$. Significantly, we discovery that the incontinuity locations of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ with $\zeta$ and $\delta$ correspond to the crossings, and the extreme points of that correspond to the anticrossings. As shown in Fig. 2(d), the green and red lines of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ are discontinuity at $\zeta=-0.3$ where the dressed energy levels exist the crossing [Fig. 2(a) and 2(g)], and the green and cyan lines of that appear the maximal values as well as the blue and green lines of that appear the minimal values at $\zeta=-0.21$, -0.02 respectively where the dressed energy levels exist the anticrossings [Fig. 2(a) and 2(g)]. We also can see that the green, blue, and cyan lines are discontinuity at $\zeta=0.07$, 0.09, 0.22 respectively [Fig. 2(e)] corresponding to the crossings of the dressed energy levels [Fig. 2(b) and 2(h)]. The green and magenta lines behave as maximal values, and the magenta and cyan lines behave as minimal values at $\zeta=-0.03$, 0.48 respectively [Fig. 2(e)] corresponding to the anticrossings of the dressed energy levels [Fig. 2(b) and 2(h)]. Similarly, the maximal values occur at $\zeta=-0.1$, -0.05, and 0.03 in the black, green, and blue lines respectively, and the minimal values occur at $\zeta=-0.05$, 0.03 in the red and cyan lines respectively [Fig. 2(f)], which correspond to the anticrossings of the dressed energy levels [Fig. 2(c) and 2(i)]. Therefore, the second order partial derivative of the photonic Hopfield coefficients is a good physical quantity to determine the crossings and anticrossings of spectra. Comparing the spectra under different configurations, one can discover that the gap in the range $\zeta\in[-0.21,-0.02]$ for the configuration HO ST and its range [Fig. 2(a)] is respectively widened and extended by the heterogeneous coupling in multichain as shown in Fig. 2(b). For the mixed configuration HO-HE ST new gaps appear near $\zeta=0$ as shown in Fig. 2(c). About $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ in Fig. 3 (d-f), 4 (e-i), and 5 (e-i), the conclusion is similar as that of Fig. 2 (d-f), thus it is no longer stated hereinbelow. When the inter-chain coupling is TT, the spectra is shown in Fig. 3(a-c). It can be seen that the gaps near $\zeta=0$ narrow, and its ranges in $\zeta$ reduce compared with Fig. 2(a-c). In the space of $\delta$, the spectra are presented in Fig. 4 (a-d) for even chain and Fig. 5 (a-d) for odd chain. In Fig. 4 (a-d) we can see that the configuration HE ST has more crossings of and energy levels of the dressed nearly-zero-energy modes than HE TT. In mixed multichains the presence of the HO chain leads also to the decrease of the crossings and energy levels, comparing Fig. 4 (a) and 4 (c). However, the crossings at $\delta=0.3$ and 0.36 [Fig. 4 (c)] covert to anticrossing at $\delta=0.19$ [Fig. 4 (d)] because of TT coupling. The energy levels of the dressed nearly-zero-energy modes cross for even chain and avoid crossing for odd chain. Therefore, there is less crossings affecting exciton transport for odd chain [Fig. 5 (a-d)].
@@
In the next section, the steady-state dynamics with $\zeta$ and $\delta$ will be shown, in which the relation between the dressed crossings (anticrossings) and the occupation number of photons can be discussed. The changes of exciton transport currents and efficiency can be explained. We aim at revealing the mechanism of exciton transport by finding the optimal range of parameters on exciton transport and comparing the spectra and steady-state dynamics for different configurations.
@@
Figure 9: (color online) The dependence of the steady-state dynamics on the inter-chain coupling strength $\zeta$ and the detuning $\Delta$ of the driving frequency relative to the TLS without cavity under TLS drive (a-f, ab-fb), and with cavity under TLS drive (g-l, gc-lc) and under TLS+cavity (the TLS and cavity drive exist simultaneously) drive (m-r) for different configurations. The corresponding physical quantities of each row are shown at right. The zones that cavity produces significant effect on exciton transport efficiency are marked by ellipses in (g-l), the corresponding zones are marked in (ab-fb) and (gc-lc). The parameters used are the same as those in Fig. 7.
@@
Figure 10: (color online) The dependence of the steady-state transport efficiency $\eta$ on the dimerization parameter $\delta$ and the detuning $\Delta$ of the driving frequency relative to the TLS, in which the corresponding conditions are written at the top and the right of this figure. The parameters used are the same as those in Fig. 7.
@@
Figure 11: (color online) The steady-state dynamics as a function of the number $L$ of the chains without a1-g1) and with (a2-h2) cavity for different configurations, in which $p_{1,i}$, $p_{L,i}$, and $p_{\rm bulk-in}$ represent respectively the occupation number of the first row, the last row, and the bulk state denoted as ”bulk-in” containing all TLSs inside of the multichains. The parameter used is $N=10$, and the others are the same as those in Fig. 7.
@@
Figure 12: (color online) The steady-state dynamics as a function of the length $N$ of chains for different configurations and number of the chains, (a-d) without cavity under TLS drive, (e-h) with cavity under TLS drive. The parameters used are the same as those in Fig. 7.

### v4

@@ II.3 Transport dynamics
For the system that exists dissipation at each TLS in the multichains, the driving field, and the coupling with environment (reservoir), its dynamics are governed by the Lindblad quantum master equation under the Born-Markov approximations and the rotating-wave approximation lindblad1976generators . Then, the exciton transport dynamics for the Hamiltonian (11) can be investigated by the master equation
@@
where the first term in the righthand side accounts for the coherent evolution under the total Hamiltonian (11), the second term characterizes the decay of excitons due to the spontaneous emission with
$D_{{\rm d},k}[\rho]=\sum^{N}_{i=1}\gamma_{d}\mathcal{L}_{\sigma^{-}_{k,i}}[\rho]+\kappa\mathcal{L}_{\rm c}[\rho],$ (14)
where $\gamma_{d}$ is the spontaneous emission rate of the TLSs, $\kappa$ is the decay rates of the cavity photons, and $\mathcal{L}_{x}[\rho]=x\rho x^{\dagger}-\dfrac{1}{2}\left\{\rho,x^{\dagger}x\right\}$ is the Lindblad superoperators with standard form. The term $D_{{\rm R},k}[\rho(t)]$ in Eq. (13) describes the exciton exchange between chain ends and exciton drain, which is given by
@@
where $\gamma_{\rm R}$ is the exciton exchange rate, $\bar{n}_{\rm R}$ is the average exciton number in the drain. The last sites of each chain exist coupling with the drain, so $k\in\{1,2,...,L\}$. The two terms in Eq. (15) represent respectively the inflow and outflow of excitons from the last column TLSs of the multichains to the drain. In order to show the net transportation of exciton, we set $\bar{n}_{\rm R}=0$, establishing unidirectional exciton transfer from the last column of the multichains to the drain without reflection, which is reasonable as the photosynthesis process in nature.
@@ II.4 Efficiency of exciton transport
Figure 2: (color online) The energy spectra of the exciton-cavity coupling system as a function of the inter-chain coupling strength $\zeta$ for different configurations HO ST (a), HE ST (b), HO-HE ST (c), with the colormap representing the photon Hopfield coefficients by numerically diagonalizing Hamiltonian (1) in single-excitation space. (d-f) The second order partial derivative of the photon Hopfield coefficients with respect to $\zeta$, which contains the locations of discontinuity and extremum corresponding to that existing evident changes in exciton transport dynamics. (g-i) The corresponding von Neumann entropy $S$ of excitons by calculating Eq. (12), in which the colors of the lines has not means just to distinguish each other. The black (violet) vertical lines and the data texts mark the locations of cavity-dressed energy level crossings (anticrossings) (a-c) producing obvious effect on exciton transport, the discontinuities (extrema) of the second order partial derivative of the photon Hopfield coefficients (d-f), as well as the saltatorial intersections (maxima or acclive intersections) of the von Neumann entropy (g-i). The corresponding parameters are $L=2$, $N=8$, $\Delta_{\rm a}=\Delta_{\rm c}=0$ eV, $h=1$ eV, $\delta=0.5$, and $\Omega_{R}=1$ eV.
@@
To quantify the exciton transport dynamics in the 2D multichain-cavity system, the exciton transport efficiency $\eta$ is defined and derived in this subsection. Generally, the exciton transport efficiency is characterized by the ratio of the outgoing $(I^{\rm s}_{\rm o})$ and input $(I^{\rm s}_{\rm i})$ exciton currents for steady state liu2019vibration ; wei2019enhanced , that is
@@
where the superscript ”s” refers to the steady state. On the other hand, the exciton occupation number can reveal the transport mechanism and determine the currents. The exciton occupation number on the $j$th site of the $k$th chain is given by
@@
Combining Eq. (13) and Eq. (17), the time evolution of $p_{k,j}(t)$ obeys
@@
where the trace is operated throughout the Hilbert space of the total system. The lefthand side describes the rate of the change of $p_{k,j}(t)$ with time, thus the first term (the second term) of the righthand side represents the input (outgoing) exciton current at the $j$th site in the $k$th chain. According to the explanation, under the TLS and TLS+cavity drive the input exciton current of the multichains can be written respectively as
@@
$I_{\rm o}(t)={\rm Tr}[\sum^{L}_{k=1}D_{{\rm R},k}[\rho(t)]\sigma_{k,N}^{+}\sigma_{k,N}^{-}].$ (22)
Pluging the Eq. (15) into the Eq. (22), under the single-excitation assumption, the form of the outgoing exciton current is simplified as
@@
where $p_{k,N}$ is the occupation number at the last site of the chain $k$.
@@ III Spectrum of coupling system
There are a large number of studies on the spectra of the SSH Hamiltonian for the open boundary conditions sirker2014boundary ; nie2020bandgap ; wei2022cavity ; cai2019observation . The topological phases of the SSH chain depend on the odevity of TLS number in 1D chain and the qubit-qubit coupling configuration. According to the Ref. cai2019observation , the dimerization parameter $\delta>0$ ($\delta<0$) corresponds to the topological winding number $\nu=1$ ($\nu=0$), i.e., the topologically nontrivial (trivial) regime. In Ref. sirker2014boundary , the spectra show that there is single zero-energy mode in whole range of $\delta$ for odd-length chain and are two localized nearly-zero-energy modes in the topologically nontrivial regime $\delta>0$ for even-length chain. The finite gap between the two edge modes leads to the formation of the hybrid edge modes that overlap with each other, generating the polaritonic states when the edge modes are dressed by cavity due to the exciton-photon coupling. Comparing to the single chain system, the presence of the inter-chain coupling induces the spectral extension of the 2D multi-chain system, adding greatly to the complexity of the system dynamics. In the exciton-cavity coupling system, the crossings and anticrossings of the edge modes dressed by the cavity can lead to the abundant exchange of excitation between the multichains and the cavity attributing to the coherent exchanges and the LZ transitions, where the polaritons increase or decrease. The polaritons act as intermediary in exciton transport, causing that the excitons are transported among site $(k,1)~{}\rightarrow$ cavity $\rightarrow$ site $(k,N)$ via bypassing dissipative intermediate TLSs wei2022cavity . Therefore, the exciton transport can be enhanced by the polaritons at the crossings and anticrossings.
Figure 6: (color online) The dependence of the steady-state occupation numbers (a1 and b1) $p_{11}$ on site (1,1), (a2 and b2) $p_{\rm bulk}$ on bulk sites, (a3 and b3) input exciton current $I_{\rm i}$, (a4 and b4) outgoing exciton current $I_{\rm o}$, and (a5 and b5) transport efficiency $\eta$ on the inter-chain coupling strength $\zeta$ (a1-5) and the dimerization parameter $\delta$ (b1-5) for different configurations without cavity. In (b1-5) the solid and dash lines correspond respectively to $N=8$ and 9. The parameters used are $\zeta=0.1$ eV, $\xi_{{\rm T},1}=\xi_{\rm c}=0.1$ eV, $\gamma_{\rm d}=0.01$ eV, $\gamma_{\rm R}=0.1$ eV, and the others are the same as those in Fig. 2.
@@
In order to analyse the spectral characteristics of the multichain-cavity coupled system, we numerically diagonalize the Hamiltonian (1) to obtain its eigenvalues. The formation of polaritons is investigated through the photon Hopfield coefficients $|\alpha_{c}|^{2}$. When the $j$th normalized eigenvector of the Hamiltonian (1) is expanded as $|\psi_{j}\rangle=\Sigma_{k}^{L}\Sigma_{i}^{N}\alpha_{k,i}|k,i\rangle+\alpha_{c}|c\rangle$, $|\alpha_{k,i}|^{2}$ and $|\alpha_{c}|^{2}$ denote the Hopfield coefficient of exciton and photon, which characterize the hybridization weights of exciton and photon components, respectively. The color maps Fig. 2 (a-c), 3 (a-c) and Fig. 4 (a-d), 5 (a-d) show respectively the spectra of the exciton-cavity coupled system as functions of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$ for different coupling configurations, and contain the photon Hopfield coefficient $|\alpha_{c}|^{2}$. One can see that the zero-energy modes exist in the spectra for all situations. $|\alpha_{c}|^{2}\neq 0$ means that the energy level is dressed by the cavity mode. It is obvious that only the nearly-zero-energy modes are distinctly dressed by the cavity mode, generating corresponding polaritonic branches, while the high excited states are hardly dressed for the given exciton-cavity coupling strength. In addition, there are the crossings (anticrossings) in the spectra, in which the crossings (anticrossings) between the nearly-zero-energy modes dressed by the cavity mode are our focus. It can be seen that the photonic content distribution on the upper and lower polaritonic branches varies significantly with $\zeta$ and $\delta$. It means that the coherent exchange of excitation or the LZ transition of photons occurs, inducing the changes of the exciton transport efficiency and currents, which will be analysed in the next section by the steady-state dynamics. Further, it can be discovered that the number and locations of crossings (anticrossings) affecting transport depend on the coupling configurations of TLSs in the range of $\zeta$ and $\delta$. In comparison, the effect of $\zeta$ on the crossings (anticrossings) is more significant than that of $\delta$.
@@
To systematically analyze the photon Hopfield coefficient $|\alpha_{c}|^{2}$, we plot the scatter maps of the second order partial derivatives $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ in Fig. 2 (d-f) and 3 (d-f), as well as $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ in Fig. 4 (e-i) and 5 (e-i), which correlate with the obvious changes of the steady-state exciton transport dynamics. To determine and distinguish accurately the locations of the crossings and anticrossings in spectra, the von Neumann entropy $S$ of excitons is calculated through Eq. (12), and is shown in Fig. 2 (g-i), 3 (g-i), 4 (i-l), and 5 (i-l). The dressed nearly-zero-energy crossings (anticrossings) identified by $S$ are marked in the spectra, which correspond to the saltatorial intersections (maxima or acclive intersections) of $S$. Significantly, we discover that the discontinuous locations of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ with $\zeta$ and $\delta$ correspond to the dressed nearly-zero-energy crossings, and the extrema points of that correspond to the dressed nearly-zero-energy anticrossings. As shown in Fig. 2(a,d,g), the discontinuities of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ at $\zeta=-0.3$ in green and red lines align with the crossings, which correspond to the saltatorial intersections of $S$. The maxima of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ in green (at $\zeta=-0.21$) and cyan (at $\zeta=-0.02$) lines, and the minima of that in blue (at $\zeta=-0.21$) and green (at $\zeta=-0.02$) lines appear at the anticrossings, which correspond to the maxima of $S$. Similar patterns emerge throughout our spectral results. We also can see that the discontinuities in green (at $\zeta=0.07$), blue (at $\zeta=0.09$), and cyan (at $\zeta=0.22$) lines [Fig. 2(e)] match the crossings [Fig. 2(b)] at the saltatorial intersections of $S$ [Fig. 2(h)]. The maxima in green (at $\zeta=-0.03$) and magenta (at $\zeta=0.48$) lines, and the minima in magenta (at $\zeta=-0.03$) and cyan (at $\zeta=0.48$) lines [Fig. 2(e)] indicate the anticrossings [Fig. 2(b)] at the maxima of $S$ [Fig. 2(h)]. In Fig. 2(c,f,i), the maxima of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ in black (at $\zeta=-0.1$), green (at $\zeta=-0.05$), and blue (at $\zeta=0.03$) lines, and the minima of that in red (at $\zeta=-0.05$) and cyan (at $\zeta=0.03$) lines correspond to the anticrossings indicated by the maxima of $S$. Therefore, the second order partial derivatives of the photon Hopfield coefficients serve as robust indicators for identifying the spectral crossings and anticrossings. About $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ in Fig. 3 (d-f), and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ in Fig. 4 (e-i) and 5 (e-i), the conclusion is similar as that of Fig. 2, thus it is no longer stated hereinbelow.
Comparing the spectra under different configurations, one can discover that the gap in the range $\zeta\in[-0.21,-0.02]$ for the configuration HO ST and its range [Fig. 2(a)] are respectively widened and extended by the heterogeneous coupling in multichains as shown in Fig. 2(b). For the mixed configuration HO-HE ST the new gaps near $\zeta=0$ appear as shown in Fig. 2(c). When the inter-chain coupling is TT, the spectra are shown in Fig. 3(a-c). It can be seen that the gaps near $\zeta=0$ narrow, and their ranges in $\zeta$ reduce compared with Fig. 2(a-c). In the space of $\delta$, the spectra are presented in Fig. 4 (a-d) for even chain and in Fig. 5 (a-d) for odd chain. In Fig. 4 (a-d) we can see that the configuration HE ST has more crossings of the dressed nearly-zero-energy modes than HE TT. What’s more, the crossings at $\delta=0.3$ and 0.36 [Fig. 4 (c)] turn into anticrossing at $\delta=0.19$ [Fig. 4 (d)] because of TT coupling. In the mixed multichains, the presence of the HO chain leads also to the decrease of the crossings, comparing Fig. 4 (a) and 4 (c). Comparing Fig. 4 with Fig. 5, one can discover that the dressed nearly-zero-energy levels cross for even-length chain and avoid crossing for odd-length chain. Therefore, we can suppose that the coherent exchange is main mechanism affecting exciton transport for former, while the LZ transition is for latter.
@@
In the next section, the steady-state dynamics with $\zeta$ and $\delta$ will be shown, in which the relation between the dressed nearly-zero-energy crossings (anticrossings) and the occupation number of photons are discussed. The exciton transport mechanism are revealed. The changes of exciton transport currents and efficiency with $\zeta$ and $\delta$ are explained. The optimal ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$ for exciton transport are found out by comparing the spectra and steady-state dynamics for different configurations.
@@
Figure 9: (color online) The dependence of the steady-state dynamics on the inter-chain coupling strength $\zeta$ and the detuning $\Delta$ without cavity under TLS drive (a-f, ab-fb), and with cavity under TLS drive (g-l, gc-lc) and under TLS+cavity drive (m-r) for HO ST (the 1st column), HO TT (the 2ed column), $\cdots$, and HO-HE TT (the 6th column). The physical quantities of each row are written at right. The zones that the cavity induces significant effect on transport efficiency $\eta$ are marked by ellipses in (g-l), the corresponding zones are marked in (ab-fb) and (gc-lc). The parameters used are the same as those in Fig. 7.
@@
Figure 10: (color online) The dependence of the steady-state transport efficiency $\eta$ on the dimerization parameter $\delta$ and the detuning $\Delta$, in which the conditions of the column and row are written at the top and right of this figure respectively. The parameters used are the same as those in Fig. 7.
@@
Figure 11: (color online) The steady-state dynamics as a function of the chain number $L$ without (a1-g1) and with (a2-h2) cavity for different configurations, in which $p_{1,i}$, $p_{L,i}$, and $p_{\rm bulk-in}$ represent respectively the occupation number of the first row, the last row, and the bulk state denoted as ”bulk-in” containing all TLSs inside of the multichains. The parameter used is $N=10$, and the others are the same as those in Fig. 7.
@@
Figure 12: (color online) The steady-state dynamics as a function of the chain length $N$ for different configurations and chain number under TLS drive, (a-d) without cavity, (e-h) with cavity. The parameters used are the same as those in Fig. 7.